i need help with this problem
Robert rented a truck for one day. There was a rental fee of $16.95, and there also was a charge of 17 cents for each mile. Robert paid $63.02 when he returned the truck. For how many miles did he drive the truck?




Answer by checkley71(8403)

16.95+.17x=63.02
.17x=63.02-16.95
.17x=46.07
x=46.07/.17
x=271 miles is the answer.
proof:
16.95+.17*271=63.02
16.95+46.07=63.02
63.02=63.02